OpenStudy (anonymous):

DOES THE FOLLOWING INFINITE SERIES CONVERGE OR DIVERGE? EXPLAIN YOUR ANSWER. (1/5)+(1/15)+(1/45)+(1/81)

OpenStudy (amistre64):

i dont see a way to compare it to something sensible as is

OpenStudy (psymon):

I can't catch on to a pattern personally. At least not yet.

OpenStudy (amistre64):

1/(5n) might be a sound comparison ....

OpenStudy (amistre64):

or 1/(5*3^n) ?

OpenStudy (psymon):

I was looking at that one before, too, but wasn't sure.

OpenStudy (amistre64):

the scant sequence provided doesnt lend much to comparison tho

OpenStudy (psymon):

It looks like some sort of p-series could work, though.

OpenStudy (psymon):

Joy :/

OpenStudy (anonymous):

\[1/5+1/15+1/45+...=1/5(1+1/3+1/3^2+...)=1/5\sum_{n=0}^{\infty} (1/3)^n\]\[=1/5 \times 1/(1-1/3)=3/10\]

OpenStudy (psymon):

The two series go different routes starting at n=3, but that's as good of a comparison to use as any.